OpenStudy (anonymous):

The base of a triangle is 8 cm greater than the height. The are is 24 cm^2. Find the height and the length of the base.

OpenStudy (anonymous):

Let the height = x So the base = x + 8 The formula for the area of a triangle is Area = (1/2)bh So, 24 = (1/2)(x)(x+8) Multiply both sides by 2, you get 48 = x(x+8) 48 = x^2 + 8x Set equal to zero because it is quadratic, x^2 + 8x - 48 = 0 Factor, (x + 12)(x - 4) = 0 x = -12 which we reject and x = 4 So the height = x = 4 cm The base = x + 8 = 12 cm

OpenStudy (anonymous):

Good Night.

OpenStudy (anonymous):

good night once again than you very much

OpenStudy (anonymous):

Welcome.